the vapour pressure of benzene at 80 oC is lowered by 10 mm by dissolving 2 g of a non- volatile substance in 78 g of benzene . the vapour pressure o f pure benzene at 80oC is 750 mm .the molecular mass of the substance will be____
 

Grade:12

2 Answers

Bhavya
7 years ago
Dear student
relative lowering of V.P (po-ps/po) = mole fraction of solute
750-740/750 = (2/M)/ 2/M + 78/78
10/750 = (2/M) / 2/M + 1
M = 148
Jms
5 years ago
po-ps/po=w2/m2 × m1/w1     
750-740/750=2/m2×78/78
   10/750=2/m2
      M2 =150.please uovote it if it is right..(question is from chapter solutions standard 12 )